Suppose the demand for good X is given by Qd = 300 – 5Px - 5Py + 20 I where Px is the price of good X. Py is the price of some other good Y, and I is income. Assume that Px is currently $1, Py is currently $2, and I is currently $50.The income elasticity is
1
Expert's answer
2020-04-27T07:49:13-0400

"Qd=300-5\u00d71-5\u00d72+20\u00d750;""Qd=" "1285;"

"E" Id"=(Qd)'"I"×\frac {I}{Qd};"

"Qd'"I"=(300-5Px-5Py+20I)'"I"=20;"

"E" Id"=20×\frac{50}{1285};"

"E" Id"=0.78;"

Answer: The income elasticity is 0.78.